# RLeXplore: Accelerating Research in Intrinsically-Motivated Reinforcement Learning

### v1

@@ 4.2 Algorithmic Baselines
In RLeXplore, we implement eight widely-recognized intrinsic reward algorithms spanning the different categories described in Section 2, namely ICM [5], RND [4], Disagreement [9], NGU [7], PseudoCounts [7], RIDE [6], RE3 [35], and E3B [8], respectively. We selected them based on the following tenet:
@@
For detailed descriptions of each method we refer the reader to Appendix A.
@@ 5 Experiments
Figure 3: Screenshots of four selected exploration games, from left to right: Montezuma’s Revenge (MR), SuperMarioBros, Procgen-Maze, and Ant-UMaze.
We design our experiments with two objectives in mind: (i) evaluating the effectiveness of our implementations and their adaptability to different RL algorithms, and (ii) investigating the tuning of intrinsic rewards to provide well-justified standard implementations.
For the first part, we use the Montezuma’s Revenge (MR) and Ant-UMaze environments, shown in Figure 3. MR is a challenging exploration task from the Atari benchmark [36], while Ant-UMaze is an exploration and locomotion task from the Gymnasium robotics benchmark [37]. We provide the experimental settings and results in Appendix E and Appendix F. In these experiments, we seamlessly integrate RLeXplore with different frameworks (CleanRL [34], RLLTE [38]) and RL algorithms (PPO [29], SAC [39]).
For the second part, we use SuperMarioBros to investigate the low-level implementation details in intrinsic reward methods that lead to more robust performance. Additionally, we use Procgen-Maze to study how to best combine intrnisic and extrinsic rewards in sparse-reward tasks.
@@ 5.1 Low-level Implementation Details of Intrinsic Rewards
The performance of intrinsic rewards is affected by various factors, which tends to vary significantly with the complexity of the task, the RL algorithm used, the architecture of the networks, algorithm-specific hyperparameters, and the joint optimization of intrinsic and extrinsic rewards. As a result, implementing and reproducing intrinsic reward algorithms is challenging. To tackle this problem, we first formulate five research questions (RQs) to investigate how various low-level implementation details impact the training of intrinsically-motivated agents. We first define an initial baseline configuration for optimizing the intrinsic rewards, shown in Table 1. These baseline settings are selected based on the most common configurations reported in the literature. Next, we address each RQ sequentially, modifying the baseline configuration for each intrinsic reward as we gather new evidence regarding their critical implementation details. This iterative process leads to the development of high-quality implementations of state-of-the-art intrinsic reward methods.
In this section, we conduct reward-free experiments (i.e., without access to extrinsic rewards) using the SuperMarioBros environment [40]. SuperMarioBros is a widely used benchmark for evaluating exploration in RL [5, 6], as efficient exploration is closely related to effectively navigating the game levels and ultimately solving the game. Additional experimental settings can be found in Appendix B.
RQ 1: The impact of observation normalization.
Figure 4: Results for RQ 1, RQ 2, RQ 3, RQ 4, and RQ 5 in SuperMarioBros-1-1. The x-axis represents the episode return normalized by the performance of the extrinsic agent after training. A normalized return of 1.0 indicates that the agent solves the 1-1 level. Each bar represents the mean and standard deviation of the normalized episode returns. Results are aggregated over 5 seeds, and each run uses 10M environment interactions.
@@
Observation normalization is crucial in deep learning to avoid numerical instabilities during optimization. Image observations, which typically range from 0 to 255, are commonly normalized to a range of 0 to 1 using Min-Max normalization by dividing by 255. However, previous studies suggest that Min-Max normalization may not be ideal for all representation learning algorithms [4].
@@
In RQ 1, we compare Min-Max normalization with using an exponential moving average (EMA) of the mean and standard deviation for observation normalization (RMS). Figure 7 indicates that using RMS for observation normalization generally reduces the variance of the agent’s performance. Importantly, some intrinsic rewards, such as RND, NGU, PseudoCounts, and RIDE, critically require RMS normalization. This results indicate that RMS normalization is crucial for intrinsic reward methods that use random networks, since the lack of normalization can result in the embeddings produced by the random networks carrying very little information about the inputs [4].
RQ 2: The impact of reward normalization.
@@
Similarly to RQ 1, reward normalization is also critical when using deep neural networks to compute the intrinsic rewards, since the scale of these rewards can be arbitrary and vary significantly over time. To mitigate the non-stationarity of intrinsic rewards, in RQ 2, we compare three normalization approaches: (1) Min-Max normalization, (2) using an RMS of the standard deviation, and (3) no reward normalization.
Surprisingly, our findings reveal that not normalizing intrinsic rewards at all can improve the performance of some intrinsic reward methods (e.g., RE3 and ICM). However, our results show that Min-Max reward normalization is a considerable option, improving the performance of the majority of the methods.
RQ 3: The co-learning dynamics of policies and representations for intrinsic rewards.
@@
Optimizing intrinsic rewards in deep RL often involves training representations in auxiliary networks (e.g., predictor network in RND, inverse dynamics encoder in ICM, forward dynamics encoders in Disagreement). However, managing the co-learning dynamics of representations and policies is challenging. In RQ 3, we explore three update strategies for representations: (1) updating representations at the same frequency as the policy, (2) updating representations 50% of the time, and (3) updating representations 10% of the time. This comparison sheds light on the trade-off between the number of gradient updates in the representations and the performance of the policy.
Our findings suggest that, in general, the learned representations of the auxiliary networks need to be caught up (i.e., equally trained) with the same data as the RL agents. Our results show that training the representations slowly (e.g., using only 10% of the data at each iteration) can cause significant performance decreases for global intrinsic rewards like ICM, Disagreement and RND. In contrast, for intrinsic rewards that utilizes episodic memories (e.g., NGU and RIDE), training the representations slowly may improve the performance to a certain extent. These observations indicate that training intrinsic rewards and RL agents does not always have to be synchronized entirely. Moreover, we can reduce the computational overhead by reducing the training time of intrinsic rewards.
RQ 4: The impact of weight initialization.
@@
Weight initialization plays a crucial role in optimizing deep neural networks, enabling faster convergence. In RQ 4, we compare two approaches for weight initialization in the auxiliary networks: (1) orthogonal weight initialization and (2) uniform weight initialization (PyTorch’s default). Note that the policy and value networks remain unchanged.
Our results show the importance of weight initialization in intrinsically-motivated RL. Specifically, we observe that using orthogonal initialization of weights leads to notable performance improvements for most intrinsic rewards, regardless of their specific optimization tasks (e.g., inverse dynamics, forward dynamics), and even in random networks (e.g., RND and RE3).
RQ 5: Is memory required to optimize intrinsic rewards?
@@
In RQ 5, we investigate whether the intrinsic rewards included in RLeXplore benefit from memory-enabled architectures. We compare the optimization of intrinsic rewards using a vanilla network and one equipped with a long-short term memory (LSTM) [41] module, while keeping PPO as the RL backbone algorithm.
@@
Some intrinsic reward methods exhibit lower performance when using LSTM policies. This observation aligns with the fact that LSTMs provide episodic context to policies, whereas most intrinsic reward methods define exploration as a global problem. Interestingly, for RIDE, which computes the state embedding changes as the intrinsic rewards, the episodic context provided by LSTMs enables agents to better optimize the intrinsic reward.
Figure 5: Overall (left) and individual (right) performance comparisons between the baselines and best configurations. For the left figure, interquartile mean (IQM), mean, median, and optimality gap (OG, lower is better) are computed over all the algorithms and runs. PoI is the probability of improvement of the best configurations compared to the baselines. Our results show that carefully-tuned implementations for the intrinsic rewards achieve significantly better exploration and task performance.
Based on the results from RQ 1-5, we have identified the optimal configurations that surpass the predefined baselines. Reward and observation normalization are notably influential on performance, and by adjusting the update proportion and using an LSTM, it is possible to further boost the performance of intrinsic rewards. Figure 5 demonstrates both the comprehensive and individual performance comparisons before and after tuning. The average probability of improvement across all intrinsic rewards is 82%, with performance increases of 86% and 66% for the aggregate IQM and mean, respectively.
@@ 5.2 Combination of Intrinsic and Extrinsic Rewards
RQ 6: Joint Optimization of Intrinsic and Extrinsic Rewards
Training agents to maximize two learning signals concurrently can be challenging. In sparse-reward environments, the objective is for agents to explore the state space by optimizing intrinsic rewards until they discover the task rewards, at which point they should focus solely on optimizing the task rewards. However, many intrinsically-motivated RL applications naively optimize the sum of intrinsic and extrinsic rewards, potentially leading to learning fuzzy value functions and suboptimal policies. In this section, we compare this approach with learning two separate value functions, one for each stream of rewards. The advantages of the latter include the ability to disentangle the effects of intrinsic and extrinsic rewards on the agent’s behavior, leading to clearer learning dynamics and potentially more efficient exploration.
For this analysis, we used the Procgen-Maze task [42] as a sparse-reward benchmark. RL agents often struggle to learn meaningful behaviors from the extrinsic reward alone in this task. We evaluate different variants of the task (e.g., 1 maze vs. 200 mazes) to examine singleton versus contextual MDPs. Figure 6 demonstrates that learning two separate value functions, which we refer as the TwoHead architecture, outperforms the naive approach of simply adding the two rewards in the complex sparse-reward environment of Procgen-Maze, both in singleton and contextual settings. Importantly, all methods outperform the extrinsic agent, especially in the 1 Maze environment.
Figure 6: (Left) During training, the extrinsic agent struggles to find the goal in the selected Maze, resulting in a reward of 0. While some intrinsic reward methods yield occasional non-zero rewards, the algorithms perform significantly better when intrinsic and extrinsic value estimation are decoupled using two distinct value heads in the agent’s network. (Right) In the Procgen variant where each maze represents a unique level, the baseline extrinsic agent achieves the goal 50% of the time, and intrinsic rewards don’t outperform the baseline significantly. We note that the presence of easier levels, where the goal may occasionally be near the agent’s starting point results in generally less sparse rewards and an easier task to learn.
@@ 5.3 Unlocking the Potential of Intrinsic Rewards
RQ 1-6 extensively discuss the tuning of intrinsic rewards under both normal and reward-free scenarios, revealing significant insights into the optimization processes. However, we aim to delve deeper into the capabilities of intrinsic rewards to address the evolving challenges in the RL community. Specifically, in RQ 7 and 8, we investigate recent developments in the exploration literature in RL, such as combined intrinsic rewards and exploration in contextual MDPs. For our experiments, we use the SuperMarioBrosRandomStages environment variant, where agents play a different level in the game at each episode. Our results indicate that the recent developments in combined intrinsic rewards merit further research, as we demonstrate that such methods can enable agents to learn exploratory behaviors of exceptional quality in both singleton and contextual MDPs.
RQ 7: Which intrinsic rewards generalize better in contextual MDPs?
In contextual MDPs, there is little shared structure across episodes, since the episodic context can vary the environment significantly. In this settings, global intrinsic rewards, which re-use experience from past episodes to compute novelty in the current episode can provide wrong estimations. Conversely, episodic intrinsic rewards, such as E3B and PseudoCounts, are specifically designed to estimate novelty within each new episode, aligning better with the dynamic nature of contextual MDPs. As shown in Figure 15, E3B achieves the highest performance among all the intrinsic rewards, while other intrinsic rewards struggle to adapt and nearly fail to learn. This distinct advantage underscores the importance of designing intrinsic rewards that are context-sensitive and capable of updating their novelty detection mechanisms based on the specific characteristics of each episode.
RQ 8: The performance of mixed intrinsic rewards.
Finally, we study the potential of combined intrinsic rewards [26]. We run experiments using all the levels in the game of SuperMarioBros, and we sample them uniformly during training. As in RQ 1-5, we do not use the extrinsic reward for training the agents but use it as an evaluation metric to show how much agents actively explore the environment.
Our results show that combined objectives enable emergent behaviors of much better quality than single objectives. Interestingly, E3B and RIDE are the best performing single objectives, and E3B+RIDE also achieves the highest performance among all the combinations. Similarly, RND and ICM, combined with other intrinsic rewards, outperform their original performance. This indicates that different intrinsic rewards can provide orthogonal gains that can be leveraged together.

### v2

@@ 4.2 Algorithmic Baselines
In RLeXplore, we implement eight widely-recognized intrinsic reward methods spanning the different categories described in Section 2, namely ICM (Pathak et al., 2017), RND (Burda et al., 2019b), Disagreement (Pathak et al., 2019), NGU (Badia et al., 2020), PseudoCounts (Badia et al., 2020), RIDE (Raileanu & Rocktäschel, 2020), RE3 (Seo et al., 2021), and E3B (Henaff et al., 2022), respectively. We selected them based on the following tenet:
@@
For detailed descriptions of each method, we refer the reader to Appendix A.
@@ 5 Experiments
Our experiments aim to achieve two main objectives: (i) highlight how intrinsic reward methods are sensitive to implementation details, and (ii) identify the best algorithmic and design choices to ensure high performance across various sparse-reward environments to demonstrate the generality and robustness of our framework.
Figure 2: Screenshots of the selected exploration games. (a) SuperMarioBros. (b) MiniGrid. (c) ALE-5. (d) Procgen-Maze. (e) Ant-UMaze.
Figure 3: Episode returns achieved by the intrinsic rewards in RLeXplore. (left) SuperMarioBros without access to the task rewards. (right) MiniGrid-DoorKey-16×16 with sparse rewards.
First, we use SuperMarioBros (SMB) without access to the environment’s rewards to study the low-level implementation details of intrinsic reward methods that drive robust exploration. We selected SMB because effective exploration within this environment strongly correlates with task performance, making it an excellent benchmark for measuring the efficacy of exploration techniques. This environment has been widely used in previous studies on exploration in RL (Pathak et al., 2019; Raileanu & Rocktäschel, 2020; Burda et al., 2019a). To further generalize our findings, we also use the MiniGrid-DoorKey-16×16 (MGD) environment, which is challenging due to the sparse rewards, making it difficult to solve with classical RL algorithms [^2] https://minigrid.farama.org/environments/minigrid/DoorKeyEnv. The effectiveness of intrinsic rewards in MiniGrid environments has also been highlighted in prior works (Raileanu & Rocktäschel, 2020; Henaff et al., 2022; 2023). With these two environments we aim to study the implementation details in both reward-free and sparse-reward tasks.
Secondly, to showcase the generalizability of RLeXplore, we evaluate our implementations in additional sparse-reward environments, including Procgen, MiniGrid, Ant-UMaze, and the set of five hard-exploration games in the arcade learning environment (ALE) suite. These experiments are designed to test how well our methods balance the use of dense intrinsic rewards with sparse extrinsic rewards across a variety of tasks. The complete set of learning curves for all the experiments is shown in Appendix E.
Lastly, we explore recent advancements in using combined intrinsic rewards (Henaff et al., 2023) to enhance exploration in contextual MDPs. Specifically, we use the full set of levels in SMB to evaluate how well both single and combined intrinsic rewards can explore various game versions and generalize their exploration across different levels.
In the following sections, we present results from SMB and MiniGrid for objective (i) and from Procgen-Maze for objective (ii). Additionally, in Appendix D, we show that using RLeXplore, we are able to reproduce and improve the performance reported in previous works for many intrinsic rewards and across multiple environments.
The design of these experiments is driven by our primary goal: to provide a general and reliable set of intrinsic reward implementations within a user-friendly framework. Instead of attempting to benchmark all methods across every possible domain, we focus on verifying the generality of each method within a carefully selected subset of popular exploration tasks. Finally, we provide the experimental details of test benchmarks and method configurations in Appendix B.
@@ 5.1 Low-level Implementation Details of Intrinsic Rewards
The performance of intrinsic rewards is affected by various factors that tend to vary with the complexity of the task. For instance, the RL algorithm used for optimization, the architecture of the networks, algorithm-specific hyperparameters, and the joint optimization of intrinsic and extrinsic rewards.
Table 2: Details of the baseline settings.
Hyperparameter Value
Observation norm. RMS
Reward norm. RMS
Weight init. Orthogonal
Update proportion 100%
with LSTM False
@@
As a result, implementing and reproducing intrinsic reward methods is challenging. To tackle this problem, we first formulate five questions to investigate how various low-level implementation details impact the training of intrinsically motivated agents. We first define an initial baseline configuration for optimizing the intrinsic rewards, shown in Table 2. These baseline settings are selected based on the most common configurations reported in the literature. Next, we address each question by modifying only one hyperparameter in the baseline configuration at a time. Finally, we evaluate the performance of these intrinsic rewards with the best parameters gathered in each question. All the experiments are conducted using SMB and MGD to investigate the effects in sparse-rewards and reward-free (i.e., without access to extrinsic rewards) scenarios, respectively.
(a) SuperMarioBros
(b) MiniGrid-DoorKey-16×16
Figure 4: Results for Q1, Q2, Q3, Q4, and Q5 in SMB (top) and MGD (bottom), which are normalized by the maximum score possibly achieved in the task. Here, ON is the observation normalization, RN is the reward normalization, UP is the update proportion, and WI is the weight initialization. The bar with a hatch mark is Combined, which refers to the results of using the best hyperparameters gathered in each question. Since RE3 only employs a fixed, randomly initialized neural network for encoding observations, there are no values in Q3. All the results are aggregated over 10 seeds, and each run uses 10M environment interactions.
Importantly, as shown in Figure 1, we keep the PPO hyperparameters fixed and the overall RL training loop unmodified throughout all the experiments in the paper in order to isolate the effect of the questions on the intrinsic reward components. Previous work has shown that PPO has many implementation details that are key to achieving great performance (Huang et al., 2022a; Engstrom et al., 2020). In the following, we study implementation details for the intrinsic reward components. The fixed PPO hyperparameters are shown in Table 8.
Q1: The impact of observation normalization.
@@
Observation normalization is crucial in deep learning to avoid numerical instabilities during optimization. Image observations, where each pixel value typically ranges from 0 to 255 per colour channel, are commonly normalized to a range of 0 to 1 using Min-Max normalization by dividing each pixel value by 255. However, previous studies suggest that Min-Max normalization may not be ideal for all representation learning algorithms (Burda et al., 2019b).
@@
In Q1, we compare Min-Max normalization with using an exponential moving average (EMA) of the mean and standard deviation for observation normalization (RMS) for the inputs to the intrinsic reward modules. RMS normalizes observations by subtracting the running mean and dividing it by the running standard deviation of all observations collected by the agent thus far. Our results shown in Figures 4 and 5 indicate that using RMS for observation normalization generally reduces the variance and achieves better asymptotic performance across all the environments of study. Importantly, some intrinsic rewards, such as RND, NGU, PseudoCounts, and RIDE, benefit significantly from RMS normalization. Critically, RND achieves zero rewards in SMB if observations are not normalized with RMS. These results indicate that RMS normalization is important for intrinsic reward methods that use random networks, since the lack of normalization can result in the embeddings produced by the random networks carrying very little information about the inputs (Burda et al., 2019b).
(a) SuperMarioBros
(b) MiniGrid-DoorKey-16×16
Figure 5: Aggregated performance of the eight intrinsic rewards with different low-level hyperparameters over 10 random seeds. The vertical dashed line represents the performance of the extrinsic agent, which only has access to the task rewards. Here, ON is the observation normalization, RN is the reward normalization, UP is the update proportion, WI is the weight initialization, IQM is the interquartile mean, OG is the optimality gap (lower is better), and Combined refers to the results of using the best hyperparameters gathered in each question. All the computation is performed using the Rliable (Agarwal et al., 2021) library.
Q2: The impact of reward normalization.
@@
Similarly to Q1, reward normalization can have a large impact when using deep neural networks to compute the intrinsic rewards, since the scale of these rewards can be arbitrary and vary significantly over time. To mitigate the non-stationarity of intrinsic rewards, in Q2, we compare three normalization approaches for the reward outputs of the intrinsic reward modules: (1) Min-Max normalization, (2) using an RMS of the standard deviation, and (3) no reward normalization.
Reward normalization smooths the optimization process, which can be beneficial for stability but can lead to slower convergence (Burda et al., 2019b). Our findings show that almost all intrinsic rewards critically require some form of reward normalization, as agents fail to explore without normalized rewards. Importantly, the latter applies to all the environments that we experiment with. Additionally, while RMS is generally the default strategy for reward normalization, our results in Figure 5 show that Min-Max normalization is a more robust option in SMB, improving the performance and reducing the variance of the majority of the methods.
Q3: The co-learning dynamics of policies and auxiliary tasks for intrinsic rewards.
@@
Optimizing intrinsic rewards in deep RL often involves training additional networks for auxiliary tasks (e.g., predictor network in RND, inverse dynamics encoder in ICM, forward dynamics encoders in Disagreement). However, managing the co-learning dynamics of the auxiliary networks and policies is challenging. In Q3, we explore three update strategies for the auxiliary networks in the intrinsic reward modules: (1) updating them at the same frequency as the policy, (2) updating them $50\%$ of the time, (3) updating them $10\%$ of the time, and (4) updating them $1\%$ of the time. This comparison sheds light on the trade-off between the number of gradient updates in the auxiliary networks and the performance of the policy. Additionally, lower update frequencies have the benefit of reducing computational overhead and training time by limiting the number of gradient updates required.
Our findings indicate that the auxiliary networks generally perform robustly across the range of studied update frequencies. Additionally, there is no clear configuration that seems generally better for all intrinsic rewards across environments, rendering this implementation detail worth tuning for specific environments.
Q4: The impact of weight initialization.
@@
Weight initialization plays a crucial role in optimizing deep neural networks, enabling faster convergence. In Q4, we compare two approaches for weight initialization in the auxiliary networks of the intrinsic reward modules: (1) orthogonal weight initialization and (2) uniform weight initialization (PyTorch’s default). Note that again, the policy and value networks remain unchanged.
Our results highlight the importance of weight initialization in intrinsically-motivated RL. Specifically, we found that orthogonal weight initialization is beneficial for most intrinsic rewards, regardless of their specific optimization tasks (e.g., inverse dynamics, forward dynamics), and even in random networks (e.g., RND and RE3). This benefit is evidenced by reduced variance in episode returns and generally higher mean returns. This observation aligns with previous research indicating that orthogonal weight initialization can improve performance stability in deep RL agents (Huang et al., 2022a; Engstrom et al., 2020). Importantly, RND is the intrinsic reward method that shows the highest variability for this implementation detail, where orthogonal weight initialization works better in SMB but worse than uniform initialization in MGD.
Figure 6: Performance of four selected intrinsic rewards in RLeXplore on the top eight most challenging tasks of the MGD suite. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Q5: Is memory required to optimize intrinsic rewards?
@@
In Q5, we investigate whether the intrinsic rewards included in RLeXplore benefit from memory-enabled architectures. We compare the optimization of intrinsic rewards using a vanilla policy network and one equipped with a long-short-term memory (LSTM) (Hochreiter & Schmidhuber, 1997) module while keeping PPO as the RL backbone algorithm.
@@
Some intrinsic reward methods exhibit significantly lower performance when using LSTM policies. This observation aligns with the fact that LSTMs provide episodic context to policies, whereas most intrinsic reward methods define exploration as a global problem.
Finally, we use the best-performing implementation details observed from Q1-5 to experiment in the set of most challenging exploration tasks from MiniGrid. Our results in Figure 6 show that with our implementations of intrinsic rewards in RLeXplore, researchers can make progress in training RL agents in challenging tasks where vanilla RL agents are unable to learn due to the sparsity of the task rewards. In summary, by systematically addressing the implementation details, our work significantly enhances the reproducibility of intrinsic reward methods. These thoughtful design choices not only improve performance but also ensure that our implementations can be reliably reproduced and generalized across various environments.
@@ 5.2 Combination of Intrinsic and Extrinsic Rewards
Q6: Joint optimization of intrinsic and extrinsic rewards.
In sparse-reward environments, the objective is for agents to explore the state space by optimizing intrinsic rewards until they discover the task rewards, at which point they should focus solely on optimizing the task rewards. However, many intrinsically motivated RL applications naively optimize the sum of intrinsic and extrinsic rewards, potentially leading to learning fuzzy value functions and suboptimal policies (Castanyer et al., 2023). In this section, we compare this common approach with learning two separate value functions, one for each reward function (Burda et al., 2019b). The advantages of the latter include the ability to disentangle the effects of intrinsic and extrinsic rewards on the agent’s behaviour, leading to cleaner learning dynamics and potentially more efficient exploration. In these settings, both value functions are used during the advantage estimation phase of PPO. Specifically, we compute separate GAE values - one using the intrinsic value function and one using the extrinsic value function. The resulting advantages are then summed to compute the policy loss term for PPO. This separation facilitates more accurate advantage estimates for each reward type, leading to improved learning dynamics.
For this analysis, we used the Procgen-Maze task (Cobbe et al., 2020) as a sparse-reward benchmark. RL agents often struggle to learn meaningful behaviours from the extrinsic reward alone in this task. We evaluate different variants of the task (e.g., 1 maze vs. 200 mazes) to examine singleton versus contextual MDPs. We note that in our framework, we do not provide different context information to the agents for singleton versus contextual MDPs (e.g., the context ID). We refer to these frameworks to formalize the agent-environment interaction when the environment remains static throughout training (i.e., singleton - 1 maze) versus when it varies at each episode (i.e., contextual - a different maze at each episode).
Figure 7 demonstrates that learning two separate value functions (Huang et al., 2022b), which we refer to as the TwoHead architecture, outperforms the naive approach of simply adding the two rewards in the complex sparse-reward environment of Procgen-Maze, both in singleton and contextual settings. Importantly, all methods outperform the extrinsic agent, especially in the 1 Maze environment.
Figure 7: (Left) During training, the extrinsic agent struggles to find the goal in the selected Maze, resulting in a reward of 0. While some intrinsic reward methods yield occasional non-zero rewards, the methods perform significantly better when intrinsic and extrinsic value estimation are decoupled using two distinct value heads in the agent’s network. (Right) In the Procgen variant, where each maze represents a unique level, the baseline extrinsic agent achieves the goal $50\%$ of the time, and intrinsic rewards don’t outperform the baseline significantly. We note that the presence of easier levels, where the goal may occasionally be near the agent’s starting point results in generally less sparse rewards and an easier task to learn.
@@ 5.3 Unlocking the Potential of Intrinsic Rewards
Q1-6 extensively discuss the tuning of intrinsic rewards under both normal and reward-free scenarios, revealing significant insights into the optimization processes. However, we aim to delve deeper into the capabilities of intrinsic rewards to address the evolving challenges in the RL community. Specifically, in Q7, we investigate recent developments in the exploration literature in RL, such as combined intrinsic rewards and exploration in contextual MDPs. For our experiments, we use the SMB-RandomStages environment variant, where agents play a different level in the game at each episode. Our results indicate that the recent developments in combined intrinsic rewards merit further research, as we demonstrate that such methods can enable agents to learn exploratory behaviours of exceptional quality in both singleton and contextual MDPs.
Q7: The performance of mixed intrinsic rewards.
We run experiments using all the levels in the game of SMB, and we sample them uniformly during training. As in Q1-5, we do not use the extrinsic reward for training the agents but use it as an evaluation metric to show how much agents actively explore the environment.
Our results show that combined objectives enable emergent behaviours of much better quality than single objectives. Interestingly, E3B and RIDE are the best performing single objectives, and E3B+RIDE also achieves the highest performance among all the combinations. Similarly, RND and ICM, combined with other intrinsic rewards, outperform their original performance. This indicates that different intrinsic rewards can provide orthogonal gains that can be leveraged together.